### v2

# Inference-Time Text-to-Video Alignment with Diffusion Latent Beam Search

Abstract
The remarkable progress in text-to-video diffusion models enables photorealistic generations, although the contents of the generated video often include unnatural movement or deformation, reverse playback, and motionless scenes. Recently, an alignment problem has attracted huge attention, where we steer the output of diffusion models based on some quantity on the goodness of the content. Because there is a large room for improvement of perceptual quality along the frame direction, we should address which metrics we should optimize and how we can optimize them in the video generation. In this paper, we propose diffusion latent beam search with lookahead estimator, which can select a better diffusion latent to maximize a given alignment reward, at inference time. We then point out that the improvement of perceptual video quality considering the alignment to prompts requires reward calibration by weighting existing metrics. This is because when humans or vision language models evaluate outputs, many previous metrics to quantify the naturalness of video do not always correlate with evaluation. We demonstrate that our method improves the perceptual quality evaluated on the calibrated reward, VLMs, and human assessment, without model parameter update, and outputs the best generation compared to greedy search and best-of-N sampling under much more efficient computational cost. The experiments highlight that our method is beneficial to many capable generative models, and provide a practical guideline that we should prioritize the inference-time compute allocation into lookahead steps for reward estimation over search budget or denoising steps.  [^1] Website: https://sites.google.com/view/t2v-dlbs  [^2] Code: https://github.com/shim0114/T2V-Diffusion-Search
1 Introduction
The remarkable progress in text-to-video diffusion models enables photorealistic, high-resolution video generation [1, 2, 3, 4]. Many future applications are anticipated, such as creating novel games [5], movies [6], or simulators to control real-world robots [7]. However, the detailed contents of the generated video often include unnatural movement or deformation, reverse playback, and motionless scenes, which should not happen in the real world. For instance, simulating factual physics in the generated video is still challenging [8, 9]. Recently, it has attracted a lot of attention to steering the output of diffusion models based on reward evaluation, quantifying the goodness of the content, which is studied as an alignment problem [10, 11]. There is a large room for improvement of perceptual quality along the frame direction in the video, and to align models with our preferences, we should address which metrics to optimize and how to optimize them.
Figure 1: (Left) Diffusion latent beam search (DLBS) seeks a better diffusion path over the reverse process; sampling $K$ latents per beam and possessing $B$ beams for the next step, which mitigates the effect from inaccurate $\operatorname{argmax}$. Lookahead (LA) estimator notably reduces the noise at latent reward evaluation by interpolating the rest of the time steps from the current latent with deterministic DDIM. (Right) DLBS achieves much better computational-efficiency than best-of-N (BoN), as achieving higher performance gains under the same execution time. LA estimator (DLBS-LA) could remarkably boost efficiency only with marginal overhead on top of DLBS.
In this paper, we propose Diffusion Latent Beam Search (DLBS) with lookahead estimator, an inference-time search over the reverse process (Figure 1; Left), which can select a better diffusion latent to maximize a given alignment reward. A lookahead estimator reduces the noise in the reward estimate, and a beam search robustly explores the latent paths, avoiding inaccurate $\operatorname{argmax}$ operation.
We then point out that the improvement of perceptual video quality considering the alignment to prompts requires reward calibration of existing metrics [12]. When we evaluate outputs by capable vision language models [13, 14] or human raters, many previous metrics to quantify the naturalness of video do not always correlate with each other, and optimal reward design to measure perceptual quality highly depends on the degree of dynamic descriptions in evaluation prompts. We design a weighted linear combination of multiple metrics, which is calibrated to perceptual quality and improves the correlation with preference.
We demonstrate that DLBS can induce high-quality outputs based on the calibrated reward, AI and human feedback (Figure 2), without model parameter update, and realize the best generation under much more efficient computational cost compared to greedy search [15, 11] and best-of-N sampling [16, 17]. The experiments also highlight that our method is beneficial to many SoTA models (e.g., Latte [18], CogVideoX [19], and Wan 2.1 [20]), and provide a practical guideline on that we should prioritize the compute allocation into lookahead steps for reward estimate more than search budget or denoising steps, at an inference time.
2 Preliminaries
Latent Diffusion Models   Latent diffusion models [21, 18] are a special class of diffusion probabilistic models [22, 23], and popular choices for high-resolution text-to-video generation [24, 25, 4], which considers the diffusion process in embedding space. Let $\textbf{x}_{0}$ be a video and encode it as $\textbf{z}_{0}=\text{Enc}(\textbf{x}_{0})$ using VAE [26]. Continuous-time forward diffusion process can be modeled as a solution to a stochastic differential equation (SDE) [27]: $d\textbf{z}=\textbf{f}(\textbf{z},t)dt+g(t)d\textbf{w},$ where $_{0}\sim p_{0}(\textbf{z})$ z ) is the latent as initial condition while $p_{t}(\textbf{z})$ is the marginal distribution of $\textbf{z}_{t}$, $\textbf{f}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, $g:\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient, and $\textbf{w}\in\mathbb{R}^{d}$ is $d$-dimensional standard Wiener process. $\textbf{f}(\cdot,\cdot)$ and $g(\cdot)$ are designed appropriately for the marginal distribution to reach $p_{T}(\textbf{z})\approx\mathcal{N}(0,\textbf{I})$ as $t\rightarrow T$ [28]. Reverse diffusion process generates samples $\textbf{z}_{0}$ through the following reverse-time SDE: $d\textbf{z}=[\textbf{f}(\textbf{z},t)-g(t)^{2}\nabla_{\textbf{z}}\log p_{t}(\textbf{z})]dt+g(t)d\bar{\textbf{w}},$ where $dt$ here is an infinitesimal negative time step from $T$ to 00 and $\bar{\textbf{w}}\in\mathbb{R}^{d}$ is a standard reverse-time Wiener process. We start this with $\textbf{z}_{T}\sim\mathcal{N}(0,\textbf{I})$. This SDE induces the marginal distribution on the data $p^{\text{pre}}(\textbf{z})$ (i.e., pre-trained diffusion models). While we omit the notation for simplicity, we consider the text-to-video generation problem, where the diffusion process is conditioned on text prompts c.
Alignment for Text-to-Video Diffusion Models In this paper, we define the alignment problem in text-to-video generation as increasing the probability of generating perceptually good video for humans, such as $\max\mathbb{E}[p(\mathcal{O}=1|\textbf{x}_{0},\textbf{c})]$ where $\mathcal{O}\in\{0,1\}$ represents if the generated video $\textbf{x}_{0}$ conditioned on c is perceptually higher quality or not. The common assumption is such a probability depends on a proxy scalar reward function $r(\textbf{x}_{0},\textbf{c})$ such as $p(\mathcal{O}=1|\textbf{x}_{0},\textbf{c})\propto\exp(\beta^{-1}r(\textbf{x}_{0},\textbf{c}))$ with $\beta\in\mathbb{R}$, and then the problem comes down to reward maximization. The proxy reward function may input the generated video $\textbf{x}_{0}$ and a prompt c.
Alignment as Stochastic Optimal Control   Previous works formulate such a reward maximization problem from the view of stochastic optimal control [29, 11, 30], where we aim to find an additional drift term $\textbf{u}(\cdot,\cdot)$ for the following reverse SDE: $d\textbf{z}=[\textbf{f}(\textbf{z},t)-g(t)^{2}\nabla_{\textbf{z}}\log p_{t}(\textbf{z})+\textbf{u}(\textbf{z},t)]dt+g(t)d\bar{\textbf{w}}$. For convenience, we adopt the change-of-variables as $\boldsymbol{\nu}_{t}:=\mathbf{z}_{T-t}$ and $\bar{\textbf{f}}(\boldsymbol{\nu},t):=\textbf{f}(\boldsymbol{\nu},t)-g(t)^{2}\nabla_{\boldsymbol{\nu}}\log p_{t}(\boldsymbol{\nu})$ because stochastic control is often based on the standard flow of time ($t:0\rightarrow T$), and then the original SDE is re-written as: $d\boldsymbol{\nu}=[\bar{\textbf{f}}(\boldsymbol{\nu},t)+\textbf{u}(\boldsymbol{\nu},t)]dt+g(t)d\textbf{w}$, where $dt$ here is an infinitesimal time step and $d\textbf{w}$ is a standard Wiener process.
Because the alignment problem comes down to reward maximization, the objective in stochastic control literature is
[…]
where $r^{\prime}(\cdot):=r(\text{Dec}(\cdot))$ evaluates the latent in the video space and $\lambda>0$. $\mathbb{E}[\cdot]$ is taken over sampling process above. In stochastic control, the optimal value function is known to be defined as,
[…]
where $p^{*}_{t}(\boldsymbol{\nu})=\frac{1}{Z}\exp(\frac{v^{*}_{t}(\boldsymbol{\nu})}{\lambda})p^{\text{pre}}_{t}(\boldsymbol{\nu})$, and obtain the optimal drift $\textbf{u}^{*}(\boldsymbol{\nu},t)=g(t)^{2}\nabla_{\boldsymbol{\nu}}\frac{v^{*}_{t}(\boldsymbol{\nu})}{\lambda}$ [31]. This optimal value function is the solution of stochastic Hamilton-Jacobi-Bellman equation [32] according to this Feynman-Kac formula [33, 34]:
$\exp\left(\frac{v^{*}_{t}(\boldsymbol{\nu})}{\lambda}\right)=\mathbb{E}_{p^{\text{pre}}}\left[\exp\left(\frac{r^{\prime}(\boldsymbol{\nu}_{T})}{\lambda}\right)|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}\right]$ (3)
and then we obtain a tractable form of the optimal drift term as:
$\textbf{u}^{*}(\boldsymbol{\nu}_{t},t)=g(t)^{2}\nabla_{\boldsymbol{\nu}}\log\mathbb{E}_{p^{\text{pre}}}\left[\exp\left(\frac{r^{\prime}(\boldsymbol{\nu}_{T})}{\lambda}\right)|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}\right].$ (4)
The intuition here is that the optimal drift pulls the current latent $\boldsymbol{\nu}$, while following the pre-trained reverse SDE, into the region achieving a higher reward at time $T$.
Figure 2: Qualitative evaluation among different search methods. We test prompts from DEVIL-high (above) and DEVIL-static (below). DLBS-LA generates more dynamic and prompt-aligned videos than GS or BoN.
3 Diffusion Latent Beam Search
We first provide a unified view of existing inference-time alignment methods through several practical approximations of optimal drift $\mathbf{u}^{*}(\boldsymbol{\nu}_{t},t)$ (Section 3.1). To mitigate errors from approximations, we propose a novel search algorithm, diffusion latent beam search with lookahead estimator (Section 3.2).
3.1 A Unified View on Practical Approximations
While Equation 4 has a relatively tractable form, it is still computationally expensive, since the expectation requires complete diffusion sampling to evaluate the latent at each time step and face numerical instability. Previous alignment methods rely on multiple-step practical approximations.
Step. 1: Jensen’s Inequality   First, when assuming $\frac{r^{\prime}(\boldsymbol{\nu}_{T})}{\lambda}$ is almost deterministic (this might hold when $t\rightarrow T$), Jensen’s inequality yields the following approximation by exchanging $\log$ and $\mathbb{E}[\cdot]$, which can be considered as a certain form of classifier guidance [35]:
$\textbf{u}^{*}(\boldsymbol{\nu}_{t},t)\approx\frac{g(t)^{2}}{\lambda}\nabla_{\boldsymbol{\nu}}\mathbb{E}_{p^{\text{pre}}}\left[r^{\prime}(\boldsymbol{\nu}_{T})|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}\right].$ (5)
Algorithm 1 Diffusion Latent Beam Search (DLBS) with Stochastic DDIM
0:  latent diffusion model $\epsilon_{\theta}$, reward function $r^{\prime}$, noise scheduling parameter $\{\alpha_{t}\}_{t=0}^{T},\{\sigma_{t}\}_{t=0}^{T}$, number of beams $B$, number of candidates $K$
1:  $\textbf{z}^{1}_{T},\cdots,\textbf{z}^{B}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$   $\triangleright$ Initial $B$ beams
2:  for $t=T$ to $1$ do
3:     for $j=1$ to $B$ do
4:         $\triangleright$ Compute the posterior mean of $\mathbf{z}_{t-1}^{j}$
5:        $\hat{\textbf{z}}_{0|t}^{j}=\frac{1}{\sqrt{\alpha_{t}}}(\textbf{z}_{t}^{j}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\textbf{z}_{t}^{j}))$
6:        $\textbf{z}^{j}_{t-1}=\sqrt{\alpha_{t-1}}\hat{\textbf{z}}_{0|t}^{j}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon_{\theta}(\textbf{z}_{t}^{j})$
7:     end for
8:     if $t>1$ then
9:        for $j=1$ to $B$ do
10:            $\triangleright$ Sample $K$ next candidate latents
11:           $\mathbf{z}_{t-1}^{ij}=\mathbf{z}^{j}_{t-1}+\sigma_{t}\epsilon_{t}^{i}$ with $\epsilon_{t}^{1},...,\epsilon_{t}^{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
12:            $\triangleright$ Estimate the clean sample from noisy latent
13:           $\hat{\mathbf{z}}_{0|t-1}^{ij}=\frac{1}{\sqrt{\alpha_{t-1}}}(\mathbf{z}_{t-1}^{ij}-\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(\mathbf{z}^{ij}_{t-1}))$
14:        end for
15:         $\triangleright$ Search $B$ higher-reward beams from $KB$ latents
16:        $\texttt{budget}:=\{(\mathbf{z}_{t-1}^{11},\hat{\mathbf{z}}_{0|t-1}^{11}),\cdots,(\mathbf{z}_{t-1}^{KB},\hat{\mathbf{z}}_{0|t-1}^{KB})\}$
17:        for $j^{\prime}=1$ to $B$ do
18:           $\mathbf{z}_{t-1}^{j^{\prime}}=\operatorname{argmax}_{\mathbf{z}^{ij}_{t-1}\in\texttt{budget}}~{}r^{\prime}(\hat{\mathbf{z}}^{ij}_{0|t-1})$
19:           $\texttt{budget}=\texttt{budget}\setminus\{(\mathbf{z}_{t-1}^{j^{\prime}},\hat{\mathbf{z}}^{\operatorname{argmax}}_{0|t-1})\}$
20:        end for
21:        $j\in\{1,\cdots,B\}\leftarrow j^{\prime}$   $\triangleright$ Reset selected $B$ indices
22:     end if
23:  end for
24:  return: $\mathbf{z}_{0}=\operatorname{argmax}_{\mathbf{z}^{j}_{0}\in\{\textbf{z}^{1}_{0},\cdots,\textbf{z}^{B}_{0}\}}~{}r^{\prime}(\mathbf{z}^{j}_{0})$
Step. 2: Tweedie’s Formula   To avoid the computationally expensive expectation, the expected reward is further approximated as $\mathbb{E}_{p^{\text{pre}}}[r^{\prime}(\boldsymbol{\nu}_{T})|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}]\approx r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t})$ where $\hat{\boldsymbol{\nu}}_{T|t}\approx\mathbb{E}_{p^{\text{pre}}}[\boldsymbol{\nu}_{T}|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}]$ is a one-step approximation of posterior mean [36], which can be calculated only with the current latent $\boldsymbol{\nu}_{t}$ without full diffusion path. Therefore, the optimal drift term can be seen as solely depending on the current time step $t$:
$\textbf{u}^{*}(\boldsymbol{\nu}_{t},t)\approx\frac{g(t)^{2}}{\lambda}\nabla_{\boldsymbol{\nu}}r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t}).$ (6)
Such a computationally tractable drift term has been leveraged for previous inference-time alignment methods via approximate guidance or twisted sequential Monte Carlo (SMC) [37]. However, as the approximated posterior mean $\hat{\boldsymbol{\nu}}_{T|t}$in intermediate steps is noisy, evaluation with the reward function for clean data $r^{\prime}(\cdot)$ may not provide a reliable signal [38]. Moreover, Equation 6 requires the reward gradient, which is not applicable to non-differentiable rewards, such as AI feedback, and is also not suitable for modalities whose reward gradient imposes a huge computational cost in practice, such as video.
Step. 3: Converting Reward Gradient into argmax   The usage of reward gradient can be converted into $\operatorname{argmax}$ operator [11, 39, 40]. The intuition here is that since the optimal drift in Equation 6 induces the diffusion latent to the direction where it maximizes the reward, we replace such a maximization with a zeroth-order search. The SDE is approximated as:
$d\boldsymbol{\nu}=\bar{\textbf{f}}(\boldsymbol{\nu},t)dt+g(t)d\textbf{w}^{*}~{}~{}\text{where}~{}~{}d\textbf{w}^{*}=\operatorname{argmax}_{d\textbf{w}}~{}r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t}).$ (7)
Note that the current diffusion latents $\boldsymbol{\nu}_{t}$ and posterior mean $\hat{\boldsymbol{\nu}}_{T|t}$ are sampled by following the standard Wiener process $d\textbf{w}$. This approximation is leveraged for inference-time alignment via greedy search [11, 39] or SMC [41] of diffusion latents. However, greedy search can result in sub-optimal generation affected by inaccurate reward estimate $r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t})$ due to its noisy input. Moreover, it can be challenging to obtain an accurate density ratio term required in SMC for a high-dimensional domain such as video generation.
3.2 Mitigating Approximation Errors via Beam Search
Existing practical algorithms based on these three approximations, such as greedy search [11, 39], fall into sub-optimal generation due to the erroneous reward evaluation with a noisy estimate of the posterior mean [36], and $\operatorname{argmax}$ operator based on them. To resolve the error accumulation, we propose a simple yet robust modification, diffusion latent beam search (DLBS) with lookahead estimator. To clearly describe the practical implementation, we use the notation of a discrete-time diffusion process in the rest of the section (see Appendix D for the continuous-time diffusion process).
Practical Implementation   We summarize the detailed sampling procedure of DLBS in Algorithm 1. For the diffusion sampler, we use stochastic DDIM [42] with a decreasing sequence $\{\alpha_{t}\}_{t=1}^{T}\in(0,1]^{T}$, noise level $\eta$, and noise schedule $\sigma_{t}=\eta\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t-1}/\alpha_{t}}$, which is equivalent to DDPM [23] when $\eta=1.0$. We initialize $B$ latent beams from the Gaussian distribution (Line 1.1), sample $K$ latents per beam in the next time step (Line 1.11), and then compute the one-step estimation of the posterior mean (Line 1.13). DLBS evaluates the estimator of posterior mean $\hat{\mathbf{z}}_{0|t-1}$ with reward function (Line 1.18) and selects Top-$B$-rewarded latent beams instead of Top-$1$ (i.e., $\operatorname{argmax}$) from $KB$ candidates (Line 1.19), which is iterated over entire reverse process from $t=T$ to $t=0$. DLBS can possess latent beams more widely than greedy search under the same budget, which mitigates error propagation due to the approximated diffusion latent evaluation.
Algorithm 2 Lookahead (LA) with Deterministic DDIM
0:  latent diffusion model $\epsilon_{\theta}$, current diffusion latent $\textbf{z}_{t-1}$, number of lookahead steps $T^{\prime}(<<T)$
1:   $\triangleright$ Run $T^{\prime}$-step deterministic DDIM starting from $\textbf{z}_{t-1}$
2:  $\tilde{t}(s)\in\{t-1,\dots,\lfloor\frac{s}{T^{\prime}}(t-1)\rfloor,\dots,\lfloor\frac{1}{T^{\prime}}(t-1)\rfloor,0\}$
3:  Select new lookahead noise schedule $\{\tilde{\alpha}_{s}\}_{s=0}^{T^{\prime}}$ for $T^{\prime}$-step interpolation of the rest of original $\{\alpha_{t^{\prime}}\}_{t^{\prime}=0}^{t-1}$
4:  $\textbf{z}_{\tilde{t}(T^{\prime})}:=\textbf{z}_{t-1}$
5:  $\tilde{\textbf{z}}_{0|\tilde{t}(T^{\prime})}=\frac{\textbf{z}_{\tilde{t}(T^{\prime})}-\sqrt{1-\tilde{\alpha}_{T^{\prime}}}\epsilon_{\theta}(\textbf{z}_{\tilde{t}(T^{\prime})})}{\sqrt{\tilde{\alpha}_{T^{\prime}}}}$
6:  for $s=T^{\prime}$ to $1$ do
7:     $\textbf{z}_{\tilde{t}(s-1)}=\sqrt{\tilde{\alpha}_{s-1}}\tilde{\textbf{z}}_{0|\tilde{t}(s)}+\sqrt{1-\tilde{\alpha}_{s-1}}\epsilon_{\theta}(\textbf{z}_{\tilde{t}(s)})$
8:     $\tilde{\textbf{z}}_{0|\tilde{t}(s-1)}=\frac{\textbf{z}_{\tilde{t}(s-1)}-\sqrt{1-\tilde{\alpha}_{s-1}}\epsilon_{\theta}(\textbf{z}_{\tilde{t}(s-1)})}{\sqrt{\tilde{\alpha}_{s-1}}}$
9:  end for
10:  return: $(\textbf{z}_{t-1},\tilde{\textbf{z}}_{0|\tilde{t}(0)})$triangleright Latent and LA estimator
Lookahead Estimator   The other source of approximation errors than the $\operatorname{argmax}$ operator is a one-step estimator of the posterior mean $\hat{\mathbf{z}}_{0|t-1}$ from Tweedie’s formula, which is still noisy, especially in earlier time steps, and leads to inaccurate reward evaluation. To reduce errors in reward evaluation, we propose a lookahead (LA) estimator $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$, which is estimated by running $T^{\prime}$-step deterministic DDIM $(1<T^{\prime}<<T)$ while equally interpolating the rest of time steps from the current latent $\mathbf{z}_{t-1}$ to $\mathbf{z}_{0}$ (Algorithm 2). While requiring additional denoising steps, its cost is almost the same as naive DLBS because most computational costs come from when we decode $\textbf{z}_{0}$ (i.e., reward evaluation). Theoretically, enlarging the lookahead steps $T^{\prime}$ monotonically tightens the upper bound on the reward-approximation error (see Appendix E). Empirically, a modest horizon ($T^{\prime}=2,3,6$) delivers substantial search improvements (Figure 8; Left), but the marginal gains saturate, so pushing $T^{\prime}$ further yields little additional benefit (see Appendix L.4).
4 Calibrating Reward to Preference Feedback
Human evaluation is one of the most valuable assessments for generative models, yet gathering human feedback at scale is prohibitively costly. A practical approach to reduce the time and cost is to leverage AI feedback from VLMs [43], which has been shown to modestly align with human judgment on video quality [44, 45] (see Appendix J). In this work, we assume that the VLM evaluation works as an oracle, and we align model outputs with the preference of VLMs, which is reasonable due to their capability and the cost to be saved. Our qualitative and quantitative evaluations also confirm that the high-rated video by VLMs is generally good for us.
However, because alignment via inference-time search requires massive reward evaluation queries, we still need to build more tractable proxy rewards that do not rely on humans or external VLM APIs. The question here is what metrics for perceptual video quality can improve the feedback from VLMs. Because the criteria of videos preferred by humans are multi-objective, maximizing a single metric may lead to undesirable generation due to over-optimization. For instance, focusing exclusively on temporal consistency or frame-by-frame quality metrics can unintentionally reduce the video’s motion magnitude. In this section, we first review the possible video quality metrics (Section 4.1), evaluate the Pearson correlation between these and VLM feedback score, and then propose a reward calibration (Section 4.2), aiming to align the existing video rewards to VLMs by considering their weighted linear combination through the brute-force search of coefficients.
Figure 3: 2D-histogram and correlation between reward functions for perceptual video quality [12] and AI feedback from Gemini [14]. A single reward (e.g., subject consistency; blue) is often not aligned well with a preference from Gemini, which happens for all the prompt sets with different degrees of dynamics. The calibrated reward, a linear combination of perceptual metrics via brute-force search (green), achieves the best Pearson correlation coefficient in all settings (statistically significant with $p<0.01$).
4.1 Metric Reward for Perceptual Video Quality
Following Huang et al. [12], we select six base reward functions for perceptual video quality (see Appendix C):
  • Subject Consistency quantifies how consistently the subject appears across video frames with DINO [46].
  • Motion Smoothness leverages the motion prior in AMT [47] to evaluate whether the generated video’s motion is smooth and physically plausible.
  • Dynamic Degree quantifies the overall magnitude of dynamic object movement by estimating optical flow [48] for each pair of consecutive frames.
  • Aesthetic Quality measures compositional rules, color harmony, and the overall artistic merit of each video frame with LAION aesthetic predictor [49].
  • Imaging Quality assesses low-level distortions (e.g., over-exposure, noise, blur) in each frame with MUSIQ predictor [50].
  • Text-Video Consistency captures how closely the content in a video aligns with a prompt with ViCLIP [51].
Figure 4: The coefficient of calibrated reward $w_{i}$ with feedback from Gemini. Each set of prompts depicting different degree of dynamics requires different mixture of perceptual video qualities.
Reward Calibration  To reflect the multi-dimensional aspect of preferred videos, we model the calibrated reward function $r^{*}(\cdot,\cdot)$ as a weighted linear combination of video quality metrics: $r^{*}(\mathbf{x}_{0},\mathbf{c}):=\sum_{i=1}^{M}w_{i}r_{i}(\mathbf{x}_{0},\mathbf{c})/\sum_{i=1}^{M}w_{i}$. The coefficient $w_{i}$ is determined by maximizing the Pearson correlation with preference feedback. We heuristically conduct a brute-force search within a reasonable range (Section 4.2).
Experimental Setup  We leverage Gemini-1.5 [14] and GPT-4o [13] as automated raters for generated videos. We provide a prompt and generated video as inputs, instructing VLMs to assign discrete scores (from 1 to 10) based on overall visual quality (e.g., clarity, resolution, brightness, and aesthetic appeal), the appropriateness of motion for either static or dynamic scenes, the smoothness and consistency of shapes and motions, and the degree of alignment with a prompt (see Appendix H).
We select four prompt sets from two distinct datasets (see Appendix F). DEVIL [52] classifies its prompts into five categories depending on the degree of movement, each further divided by subject type (e.g., cat, horse, truck, nature, etc.). We focus on three of five categories (high, medium, and static) and select one prompt randomly from each subject-subdivision within a chosen category. We also draw 30 random captions from the test split of MSRVTT [53], widely used as a video benchmark.
We generate 64 videos per prompt from pre-trained Latte [18] using the DDIM sampler with $T=50$ and $\eta=0.0$ to examine the correlation among AI feedback and perceptual quality metrics. We also prepare candidates for the calibrated reward by choosing the combination of weights $w_{i}\in\{0.0,0.25,0.5,0.75,1.0\}$ and use those later to rank them based on the correlation with AI feedback.
4.2 Correlation and Reward Calibration
Figure 3 illustrates the 2D-histogram and corresponding correlation between each metric and feedback from Gemini. Relying on a single metric often yields low correlation, which supports the multifaceted nature of perceptual video quality. See Appendix I.1 for further results, where we can see that the relative importance of each metric depends on the movement in the prompts; in highly dynamic DEVIL-high, the dynamic degree correlates more strongly with VLMs than consistency metrics. Conversely, subject consistency and motion smoothness play more prominent roles in less-dynamic DEVIL-medium or DEVIL-static. Because the aesthetic score focuses on frame-by-frame visual quality, it tends to correlate strongly with VLM in low-motion scenarios. Rapid movements and frequent transitions often introduce motion blur or abrupt changes in composition, reducing the frame-level aesthetic quality and thus weakening its correlation with VLMs.
Reward calibration, a weighted linear combination of these metrics, yields the highest correlation with Gemini (Figure 3, green). We select the best coefficients among brute-force candidates, based on the correlation with Gemini, for each set of prompts that have a different description of the degree of movement (Figure 4). The dynamic degree and motion smoothness exhibit pronounced variation among DEVIL-high, -medium, and -static, and the weights of other factors, such as text-video consistency, subject consistency, and aesthetic considerations, are also adjusted accordingly. In addition, Appendix K presents results from best-of-64 sampling with a single metric or calibrated reward, where a single metric often leads to over-optimization. This highlights the importance of reward calibration, appropriately weighting multiple criteria, as aligning with a request in the prompt.
Figure 5: (Left) Comparison among diffusion latent beam search (DLBS), best-of-N (BoN), and greedy search (GS). We measure the performance in terms of a combinational reward calibrated to Gemini. DLBS improves all the calibrated rewards the best as the search budget $KB$ increases (especially $KB=16,32$), while BoN and GS, in some cases, eventually slow down or saturate the performance. Notably, an LA estimator with a small search budget ($KB=8,T^{\prime}=6$) is comparable to or even outperforms DLBS ($KB=32$). (Right) Optimal balance between the number of latent $K$ and the number of beams $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we peak around $K=4,8,16$, which is about 25–50% of the budget.
5 Inference-Time Text-to-Video Alignment
Experimental Setup  We use the same prompts and Gemini-/GPT-calibrated rewards as in Section 4. We compare the following inference-time search methods with a noise level $\eta=1.0$ for DDIM:
  • Best-of-N Sampling (BoN):  We initialize $B$ latents and they follow the reverse process independently ($K=1$). At $t=0$, we evaluate the reward and select the best.
  • Greedy Search (GS):  At each denoising step, we select the best-rewarded diffusion latent ($B=1$) from $K$ candidates sampled in a reverse process.
  • DLBS:  Given the budget $KB$, we sweep possible combinations in terms of power of 2 (e.g., $K=8,B=2$), and report the best results except for the case with $K=1$ and $B=1$.
  • DLBS-LA:  We combine DLBS with a lookahead estimator from the 6-step deterministic DDIM.
We normalize the rewards with min/max values so that the range is around $[0,1]$. Our experiments aim to assess: (1) scaling the search budget and computational costs for efficient resource allocation (Section 5.1); (2) evaluating alignment performance with feedback from humans and VLMs (Section 5.2); (3) assessing scalability to capable SoTA models (Section 5.3); (4) performing detailed ablations on LA steps $T^{\prime}$, and robustness to diverse and complex prompts (Section 5.4); (5) quantitative analysis on the diversity of generated video (Appendix L.7).
5.1 Scaling Search Budget and Computational Cost
Figure 5 (Left) measures the combinatorial reward calibrated to Gemini while increasing the search budget $KB\in\{1,2,4,8,16,32\}$. DLBS improves all the calibrated rewards the best as $KB$ increases (especially $KB=16,32$), while BoN and GS eventually slow down or saturate the performance in some cases. See Appendix L.1 for results with GPT calibrated reward and Appendix L.2 for the results with $KB=64$, where we still observe the improvement.
Figure 5 (Right) demonstrates the scaling trend of DLBS, proportional to the search budget, under various choices of $K$. The results show that there is an optimal balance between the number of latent $K$ and the number of beams $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we peak around $K=4,8,16$, which is about 25–50% of the budget. This implies that balancing possession and exploration of diffusion latents in DLBS helps search for the best outputs robustly. See Appendix L.3 for further results.
We also analyze how alignment performance scales with different DLBS configurations under fixed computational budgets, using function evaluations (NFE) and wall-clock time as cost measures. As shown in Figure 6 (Left), DLBS consistently outperforms BoN and GS across all budgets, demonstrating superior efficiency in utilizing compute. Adding the LA estimator further amplifies this advantage, offering substantial performance gains with minimal overhead. In contrast, increasing the number of diffusion steps $T$ (DLBS-T; $KB=8$, $T\in\{100,200\}$) results in only marginal improvements despite the higher computational cost. Our results suggest a clear strategy for inference-time budget allocation: prioritize enabling the LA estimator and increasing the search budget $KB$ leads to substantial performance gains, while increasing the diffusion steps $T$ provides limited benefit relative to its computational cost.
Figure 6: (Left) DLBS achieves alignment performance gains more efficiently than BoN and GS under the same number of function evaluations (NFE) or execution time. Increasing the search budget provides larger improvements under an equivalent computational cost than scaling the number of diffusion steps (DLBS-T; $KB=8,T=100,200$). Employing the LA estimator (DLBS-LA) further amplifies these gains, only with marginal overhead, yielding remarkably better efficiency than BoN or GS. (Right) DLBS and DLBS-LA help the latest SoTA models, CogVideoX-5B [19] and Wan 2.1-14B [20], improve the generated video quality.
5.2 Evaluation with AI and Human Feedback
As discussed in Section 4, we obtain a manageable reward function through the reward calibration, which reduces the cost for frequent evaluation queries in inference-time search. While DLBS efficiently improves the calibrated reward (Figure 5), a natural question is whether DLBS can improve an actual assessment by VLMs or humans by optimizing their calibrated rewards. We first use each calibrated reward for DLBS, then evaluate the quality using discrete scores (from 1 to 10) from Gemini or GPT-4o. Figure 7 (Left) demonstrates that DLBS maximizing calibrated rewards can improve the original preference feedback from VLMs, as we grow the search budget.
Next, we conduct an evaluation with VideoScore [54], a metric trained on human judgments that evaluates videos across five quality criteria. Figure 7 (Middle) shows that DLBS significantly improves the quantitative evaluation based on human evaluation. Lastly, we perform pairwise comparisons between DLBS-LA and GS ($KB=1$) by three human evaluators (Figure 7; Right). The results confirm that, whatever models or prompts we choose, the quality of content generated by DLBS-LA consistently outperforms that of a baseline. This emphasizes that our proposed method, integrating reward calibration and beam search in a latent space, effectively enhances perceptual video quality. See Appendix L.10 for the details.
Figure 7: (Left) DLBS on calibrated reward can improve the original preference feedback from VLMs. As in Figure 5, we use each calibrated reward (Figure 4) for DLBS and evaluate the quality with Gemini or GPT-4o. (Middle) DLBS on calibrated reward also improves another qualitative metric, the most, VideoScore [54], which is not involved in a reward calibration. (Right) The pairwise comparisons between DLBS-LA and GS ($KB=1$) by three human raters, when DLBS-LA searches for the best latents in a diffusion process.
5.3 Scaling Model Parameters and Capabilities
We scale up the base diffusion model from Latte to the latest SoTA models such as CogVideoX-5B [19] and Wan 2.1-14B [20], and evaluate if DLBS can improve the generation quality from those larger models. Since base models are more capable here, we adopt more challenging prompts from DEVIL-very-high (22 prompts) and Movie Gen Video Bench [55] (20 prompts). See Appendix F and I.2 for the details. Figure 6 (Right) shows that our methods achieve significant improvement in calibrated reward for both models. As shown in Figure 7 (Right), human evaluation also supports that our proposed method could generally work well with any text-to-video models, even with more capable models in the future. See Appendix L.9 for other results with CogVideoX and Wan 2.1.
Figure 8: (Left) Comparison among different LA steps $T^{\prime}$ on MSRVTT-test ($KB=8$). The performance improves as the number of LA steps increases. (Right) Perceptual quality comparison on a large and complex prompt set (1,003 prompts from Movie Gen Video Bench [55]). DLBS-LA can generalize to diverse prompts.
5.4 Ablation Study
Lookahead Steps for Reward Estimate  We scale up LA steps $T^{\prime}$ to obtain an accurate reward estimate. We use MSRVTT-test ($KB=8$) and Gemini reward for experiments. Figure 8 (Left) shows that as the number of LA steps increases, the performance improves more. Even $T^{\prime}=2,3$ significantly outperforms the posterior mean, which is often used in prior works [15, 11, 41]. This is because the sub-optimal performance of inference-time search comes from the approximation errors and LA estimator can notably reduce them. As shown in Figure 5, DLBS-LA $(KB=8,T^{\prime}=6)$ achieves comparable or even outperforming results with DLBS $(KB=32)$. It is quite beneficial to spend a computation to estimate the reward accurately. See Appendix L.4 for further discussions. In addition, an ablation study on diffusion steps is shown in Appendix L.5.
A Large and Complex Prompt Set  To confirm the robustness of our approach on a large and highly diverse prompt set, we compare GS ($KB{=}1$) with DLBS-LA ($KB{=}8,\;T^{\prime}{=}6$) with all the prompts in Movie Gen Video Bench [55], which comprises 1,003 prompts (Figure 8; Right). We observe that DLBS-LA consistently delivered substantially higher alignment rewards, demonstrating that DLBS generalizes effectively to complex prompt distributions.
6 Related Works
Classifier guidance [35, 56] has been the most popular to enhance text-content alignment. On top of that, recent works [11, 39] leverage reward or external feedback at inference time by selecting better latents [57], which probably achieve higher rewards during the reverse process. Kim et al. [58] propose twisted SMC [37] with reward gradient, which is not suitable for non-differentiable feedback and for domains such as video, where reward gradient needs a huge memory cost. Gradient-free methods [59, 17] such as SMC [41] or greedy search [15] often exhibit sub-optimal results affected by inaccurate reward estimates from noisy latents. Yeh et al. [60] uses ODE to estimate the reward, but it highly depends on Karras sampler [28] to avoid numerical instability. In contrast, we address the error propagation from inaccurate reward estimates with beam search and lookahead estimator via deterministic DDIM, which is more popular and stable. Our methods work more scalably when allocating more computation budget at inference time. See Appendix M for further related works.
7 Discussion and Limitation
Our reward calibration assumes that VLMs work as a proxy of human evaluation, and we demonstrate both qualitative and quantitative improvement of the video quality. It would be important to extend it to human feedback, not limited to VLMs. In practice, we often do implicit or explicit best-of-N sampling for video generation. In contrast, DLBS-LA exhibits much better computational efficiency. Spending more computation at inference time significantly improves perceptual quality, but it is orthogonal compared to speeding up the sampling process via distillation [61, 62], architecture changes [63], or parallel sampling [64]. We believe both high-quality and speedy sampling have practical needs and should be balanced.
8 Conclusion
This paper studies which metrics we should optimize and how to optimize them for better text-to-video generation. We point out that feedback from humans or capable VLMs reflects multiple dimensions of video quality, so optimizing an existing metric alone is insufficient; rather, we should calibrate the reward by combining. Our DLBS with LA estimator reduces the error propagation from the inaccurate reward estimate. We demonstrate DLBS is the most scalable, efficient and robust inference-time search that significantly improves video quality under the same computational costs. We hope our work encourages more uses of inference-time computation for text-to-video models.
Acknowledgements
We thank Po-Hung Yeh, Shohei Taniguchi, Kuang-Huei Lee, Arnaud Doucet, Heiga Zen, and Yusuke Iwasawa for their support and helpful discussion on the initial idea of this work. HF was supported by JSPS KAKENHI Grant Number JP22J21582 (by March, 2025), and MS was supported by JSPS KAKENHI Grant Number JP23H04974.
Appendix A Broader Impacts
Our work contributes to the progress of text-to-video by focusing on improving the perceptual quality and fidelity of generated videos, specifically addressing issues like unnatural movement, deformation, and temporal inconsistencies, through the inference-time alignment algorithm. Such advancements hold immense potential for revolutionizing creative fields and enabling new applications in gaming, filmmaking, and robotics.
On the other hand, the ability to generate highly realistic videos raises concerns about the potential for misuse in creating deceptive content, including deepfakes and misinformation. Like other generative models, such as large language models, text-to-video models, and their inference-time search, may inherit and amplify biases present in the training data due to the misalignment. This might lead to the generation of videos that perpetuate harmful stereotypes or underrepresent certain groups.
Lastly, by focusing on inference-time alignment, our method promotes more use of computational resources at test time. On one side, this may increase the environmental footprint for running large generative models and on the other side, our detailed recipe can contribute to designing efficient use of resources and reducing the footprint associated with training. We believe that discussing this aspect is crucial as the scale of these models continues to grow.
Appendix B Implementation Details
B.1 Code.
Our implementation for the experiment is available at https://anonymous.4open.science/r/T2V-Diffusion-Search-537B.
B.2 Models.
Our experiments cover three text-to-video diffusion models.
  • Latte [18]: a T5-conditioned Latent Diffusion Transformer with 1.1 B parameters, built on PixArt-$\alpha$ [65, 66].
  • CogVideoX [19]: a larger DiT-based model with 2B or 5B parameters. We mainly used CogVideoX-5B.
  • Wan 2.1 [20]: a DiT-based flow model with 1.3B or 14B parameters. We use Wan 2.1-1.3B for reward calibration and Wan 2.1-14B for inference-time search experiments.
B.3 Hyperparameters
  • Latte: DDIM scheduler with a linear noise schedule $(\beta_{\text{start}}=1.0\times 10^{-4},\ \beta_{\text{end}}=2.0\times 10^{-2})$ and classifier-free guidance scale $w_{\text{cfg}}=7.5$.
  • CogVideoX: DDIM scheduler with the original settings and $w_{\text{cfg}}=6.0$. Owing to computational resource constraints, we limit the frame length to 17 per sample.
  • Wan 2.1: DPMSolver++ with guidance scale $w_{\text{cfg}}=5.0$. For the same computational reasons, the spatial resolution is limited to $832\times 480$ and the frame length to 33 per sample.
B.4 Hardware configuration
  • Latte: FP16 inference on a single NVIDIA A100 (40 GB), batch size 1.
  • CogVideoX: BF16 inference on a single NVIDIA A100 (40 GB), batch size 1.
  • Wan 2.1: FP16 inference on four NVIDIA H100s (80 GB each), batch size 1.
B.5 AI-feedback endpoints
We use API endpoints: gemini-1.5-pro-002 and gpt-4o-2024-11-20.
Appendix C Details of Metric Rewards
Subject Consistency  We adopt the subject consistency metric proposed in VBench [12] to quantify how consistently a subject is depicted across consecutive video frames. Concretely, for each frame $i$ in a video, we extract a feature representation $\mathbf{d}_{i}$ using DINO [46] with a ViT-B/16 [67] backbone. Let $\langle\mathbf{d}_{i},\mathbf{d}_{j}\rangle$ denote the cosine similarity between the features $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$. Then, VBench defines the subject consistency metric as follows:
$R_{\text{subject}}=\frac{1}{T-1}\sum_{t=2}^{T}\frac{1}{2}\bigl{(}\langle\mathbf{d}_{1},\,\mathbf{d}_{t}\rangle+\langle\mathbf{d}_{t-1},\,\mathbf{d}_{t}\rangle\bigr{)}.$ (8)
DINO, which is trained in a self-supervised manner using unlabeled images and image augmentations, does not explicitly suppress intra-class variations. As a result, it remains particularly sensitive to identity shifts within the same subject, making it well-suited for evaluating subject consistency across frames.
Motion Smoothness  We adopt the frame-interpolation-based metric originally proposed in VBench [12] to assess whether a generated video’s motion is smooth and physically plausible. In particular, this metric leverages the motion prior from AMT [47], employing its AMT-S variant for frame reconstruction. Concretely, let $\bigl{[}\mathbf{f}_{0},\mathbf{f}_{1},\mathbf{f}_{2},\ldots,\mathbf{f}_{2n}\bigr{]}$ denote the frames of a generated video. We remove each odd-numbered frame to obtain a lower-frame-rate sequence $\bigl{[}\mathbf{f}_{0},\mathbf{f}_{2},\mathbf{f}_{4},\ldots,\mathbf{f}_{2n}\bigr{]}$, and rely on AMT-S to reconstruct the missing frames $\bigl{[}\hat{\mathbf{f}}_{1},\hat{\mathbf{f}}_{3},\ldots,\hat{\mathbf{f}}_{2n-1}\bigr{]}$. We then compute the Mean Absolute Error (MAE) between these reconstructed frames and the original odd-numbered frames, denoting this measure by $R_{\text{smoothness}}$. Finally, following the normalization scheme introduced in VBench, we define:
$R_{\text{smoothness-norm}}=\frac{255-R_{\text{smoothness}}}{255},$ (9)
which ensures that the final score lies in the range $[0,1]$, with higher values indicating smoother motion. This measure leverages the motion prior in AMT to evaluate whether the generated video’s motion is smooth and physically plausible. We remove each odd-numbered frame, then use AMT-S to reconstruct those frames based on short-term motion assumptions.
Dynamic Degree  This measure quantifies the overall magnitude of dynamic object movement. Let $T$ be the total number of frames in the generated video. For each pair of consecutive frames $t$ and $t+1$, we estimate the optical flow $\mathbf{v}_{t}$ using RAFT [48], compute its norm $\|\mathbf{v}_{t}\|$, and sum these values across all frames:
$R_{\text{dynamics}}=\sum_{t=1}^{T-1}\|\mathbf{v}_{t}\|.$ (10)
We then apply a logarithmic transformation to $R_{\text{dynamics}}$ and divide by 16:
$R_{\text{dynamics-rescaled}}=\frac{\log\bigl{(}R_{\text{dynamics}}\bigr{)}}{16}.$ (11)
This rescaling helps ensure that the value range of $R_{\text{dynamics-scaled}}$ is roughly comparable to other metrics in our evaluation.
Aesthetic Quality  This criterion evaluates compositional rules, color harmony, and overall artistic merit on a per-frame basis. Concretely, for each frame $i$ in a video, we extract a CLIP image embedding $\mathbf{c}^{image}_{i}$ using the CLIP ViT-L/14 model [68]. We then feed $\mathbf{c}^{image}_{i}$ into the LAION aesthetic predictor [49], which assigns a raw rating $r_{i}\in[0,10]$. To normalize these scores to the $[0,1]$ range, we set
$r^{\prime}_{i}=\frac{r_{i}}{10}.$ (12)
Let $T$ be the total number of frames. The final aesthetic reward is then obtained by taking the average of the normalized ratings across all frames:
$R_{\text{aesthetic}}=\frac{1}{T}\sum_{i=1}^{T}r^{\prime}_{i}.$ (13)
Because the LAION aesthetic predictor leverages CLIP embeddings instead of raw images, it captures higher-level features related to composition, color harmony, and artistic appeal.
Imaging Quality  This indicator assesses low-level distortions (e.g., over-exposure, noise, blur) in each generated frame. We adopt the MUSIQ predictor [50], trained on the SPAQ dataset [69]. The frame-wise score is normalized to [0, 1] by dividing by 100, and the final video score is the mean of these normalized values across all frames in the same way as Equation 12 and Equation 13.
Text-Video Consistency This measure captures how closely a generated video’s content aligns with its text prompt. We employ ViCLIP [51], a model pre-trained on a 10M video-text dataset and fine-tuned to handle temporal relationships, to embed both the video frames and the text. Since ViCLIP computes embeddings from 8-frame inputs, we sample 8 frames from each video. Let $\mathbf{v}^{\text{video}}$ denote the resulting video embedding and $\mathbf{v}^{\text{text}}$ denote the text embedding. We then define the final alignment score as the cosine similarity between these embeddings:
$R_{\text{tv-consistency}}=\langle\mathbf{v}^{\text{video}},\mathbf{v}^{\text{text}}\rangle$ (14)
Appendix D Algorithms with Continuous-time Diffusion Process
In this section, we present our algorithms for a continuous-time diffusion process. For Wan 2.1 [20], we integrated the proposed search algorithm into DPMSolver++ [70], a widely used continuous-time solver. Algorithms 3 and 4 present the pseudocode. Although we present the first-order variant for clarity, the procedure extends straightforwardly to higher-order formulations. Throughout this section, we adopt the notation of Lu et al. [70].
Algorithm 3 Diffusion Latent Beam Search (DLBS) with SDE-DPMSolver++
0:  signal prediction latent diffusion model $z_{\theta}$, reward function $r^{\prime}$, time steps $\{t_{s}\}_{s=0}^{M}$, noise scheduling parameter $\{\alpha_{t_{s}}\}_{s=0}^{M},\{\sigma_{t_{s}}\}_{s=0}^{M}$, number of beams $B$, number of candidates $K$
1:  $\textbf{z}^{1}_{t_{0}},\cdots,\textbf{z}^{B}_{t_{0}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$   $\triangleright$ Initial $B$ beams
2:  for $s=1$ to $M$ do
3:     for $j=1$ to $B$ do
4:         $\triangleright$ Update one step to produce $\mathbf{z}_{t_{s}}^{j}$
5:        $\textbf{z}^{j}_{t_{s}}=\frac{\sigma_{t_{s}}}{\sigma_{t_{s-1}}}e^{-h}\textbf{z}_{t_{s-1}}^{j}+\alpha_{t_{s}}(1-e^{-2h})z_{\theta}(\textbf{z}_{t_{s-1}}^{j})$
6:     end for
7:     if $s<M$ then
8:        for $j=1$ to $B$ do
9:            $\triangleright$ Sample $K$ next candidate latents
10:           $\mathbf{z}_{t_{s}}^{ij}=\mathbf{z}^{j}_{t_{s}}+\sigma_{t}\sqrt{e^{-2h}-1}\epsilon_{t_{s-1}}^{i}$ with $\epsilon_{t_{s-1}}^{1},...,\epsilon_{t_{s-1}}^{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
11:            $\triangleright$ Estimate the clean sample from noisy latent
12:           $\hat{\mathbf{z}}_{t_{M}|t_{s}}^{ij}=\frac{\sigma_{t_{M}}}{\sigma_{t_{s}}}\mathbf{z}_{t_{s}}^{ij}-\alpha_{t_{M}}(e^{-h}-1)z_{\theta}(\mathbf{z}^{ij}_{t_{s}}))$
13:        end for
14:         $\triangleright$ Search $B$ higher-reward beams from $KB$ latents
15:        $\texttt{budget}:=\{(\mathbf{z}_{t_{s}}^{11},\hat{\mathbf{z}}_{t_{M}|t_{s}}^{11}),\cdots,(\mathbf{z}_{t_{s}}^{KB},\hat{\mathbf{z}}_{t_{M}|t_{s}}^{KB})\}$
16:        for $j^{\prime}=1$ to $B$ do
17:           $\mathbf{z}_{t_{s}}^{j^{\prime}}=\operatorname{argmax}_{\mathbf{z}^{ij}_{t_{s}}\in\texttt{budget}}~{}r^{\prime}(\hat{\mathbf{z}}^{ij}_{t_{M}|t_{s}})$
18:           $\texttt{budget}=\texttt{budget}\setminus\{(\mathbf{z}_{t_{s}}^{j^{\prime}},\hat{\mathbf{z}}^{\operatorname{argmax}}_{t_{M}|t_{s}})\}$
19:        end for
20:        $j\in\{1,\cdots,B\}\leftarrow j^{\prime}$   $\triangleright$ Reset selected $B$ indices
21:     end if
22:  end for
23:  return: $\mathbf{z}_{t_{M}}=\operatorname{argmax}_{\mathbf{z}^{j}_{t_{M}}\in\{\textbf{z}^{1}_{t_{M}},\cdots,\textbf{z}^{B}_{t_{M}}\}}~{}r^{\prime}(\mathbf{z}^{j}_{t_{M}})$
Algorithm 4 Lookahead (LA) with DPMSolver++
0:  signal prediction latent diffusion model $z_{\theta}$, current diffusion latent $\textbf{z}_{t_{s}}$, number of lookahead steps $M^{\prime}(<<M)$
1:   $\triangleright$ Run $M^{\prime}$-step deterministic DPMSolver++ starting from $\textbf{z}_{t_{s}}$
2:  $\tilde{s}(u)\in\{s,\dots,\lfloor\frac{M^{\prime}-u}{M^{\prime}}s+\frac{u}{M^{\prime}}M\rfloor,\dots,\lfloor\frac{1}{M^{\prime}}s+\frac{M^{\prime}-1}{M^{\prime}}M\rfloor,M\}$
3:  Select new lookahead noise schedule $\{\tilde{\alpha}_{t_{u}}\}_{u=0}^{M^{\prime}}$ for $M^{\prime}$-step interpolation of the rest of original $\{\alpha_{t_{s}^{\prime}}\}_{s^{\prime}=0}^{M}$
4:  $\textbf{z}_{t_{\tilde{s}(0)}}:=\textbf{z}_{t_{s}}$
5:  for $u=1$ to $M^{\prime}$ do
6:     $\tilde{\textbf{z}}_{t_{\tilde{s}(u)}|t_{\tilde{s}(u-1)}}=\frac{\sigma_{t_{\tilde{s}(u)}}}{\sigma_{t_{\tilde{s}(u-1)}}}\mathbf{z}_{t_{\tilde{s}(u-1)}}^{ij}-\alpha_{t_{\tilde{s}(u)}}(e^{-h}-1)z_{\theta}(\mathbf{z}^{ij}_{t_{\tilde{s}(u-1)}}))$
7:  end for
8:  return: $(\textbf{z}_{t_{s}},\tilde{\textbf{z}}_{t_{M}|t_{\tilde{s}(0)}})$triangleright Latent and LA estimator
Appendix E Theoretical Analysis on Lookahead Estimator
Consider the Lookahead estimator described in Algorithm 2, which obtains the state $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}$ by performing $T^{\prime}$ steps of DDIM (or another diffusion-based sampling) with $\eta=0.0$. Our goal is to show that, as $T^{\prime}$ grows, the reward estimate $r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})$ converges to $r^{\prime}(\mathbf{z}_{0})$, thereby improving estimation accuracy.
Let $\mathbf{z}_{t-1}$ be a state in the latent space from which we wish to recover the initial latent $\mathbf{z}_{0}$. By applying $T^{\prime}$ steps of DDIM with $\eta=0.0$, we obtain an approximation $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}$. From prior work [71], the error $\|\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}-\mathbf{z}_{0}\|$ scales as follows:
$\|\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}-\mathbf{z}_{0}\|\leq\begin{cases}\mathcal{O}(1/T^{\prime})&\text{(DDIM)},\\ \mathcal{O}(1/\sqrt{T^{\prime}})&\text{(DDPM)},\\ \mathcal{O}(1/(T^{\prime})^{n})&\text{(an $n$-th order solver)}.\end{cases}$
Hence, increasing $T^{\prime}$ yields a progressively better approximation of $\mathbf{z}_{0}$.
Assume $\mathbf{z}_{0}$ is the latent representation at time $t=0$. By the Continuous Mapping Theorem, if $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}\to\mathbf{z}_{0}$ as $T^{\prime}\to\infty$, then for any continuous function $f$, we have
$f(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})\;\to\;f(\mathbf{z}_{0}).$
Setting $f(\cdot)=r^{\prime}(\cdot)$, where $r^{\prime}$ is our reward model, yields
$r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})\;\to\;r^{\prime}(\mathbf{z}_{0}),$
as $T^{\prime}\to\infty$.
We further assume that the reward model $r^{\prime}(\cdot)$ is Lipschitz continuous with Lipschitz constant $L$. Then for any two latent states $\mathbf{z}_{a}$ and $\mathbf{z}_{b}$, the reward estimates satisfy
$|r^{\prime}(\mathbf{z}_{a})-r^{\prime}(\mathbf{z}_{b})|\;\leq\;L\|\mathbf{z}_{a}-\mathbf{z}_{b}\|.$
Hence, the order of the error in $r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})$ tracks the order of the error in $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}$ itself. Explicitly,
$\bigl{|}r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})-r^{\prime}(\mathbf{z}_{0})\bigr{|}\;\leq\;L\,\|\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}-\mathbf{z}_{0}\|,$
implying that an $\mathcal{O}(1/T^{\prime})$ (or better) approximation in latent space implies an $\mathcal{O}(1/T^{\prime})$ (or correspondingly better) approximation in the reward space.
As $T^{\prime}$ increases, $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}$ converges to $\mathbf{z}_{0}$, and consequently $r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})$ converges to $r^{\prime}(\mathbf{z}_{0}$. Because the reward model is Lipschitz continuous, this convergence ensures that the error in reward estimation decreases at the same order as the error of the latent approximation. Therefore, employing the LA estimator with a larger $T^{\prime}$ yields a more accurate reward estimate.
Appendix F List of Prompts
MSRVTT-test
  1. 1.
    a woman is singing on stage about that one person being the one she wants
  2. 2.
    someone is filming a parked car in the parking lot
  3. 3.
    a cat is feed it s babies and a rabbit
  4. 4.
    mario game with bombs
  5. 5.
    someone is browsing a set of games on their console
  6. 6.
    a game is being played
  7. 7.
    a man holds a very large stick
  8. 8.
    a yellow-haired girl is explaining about a game
  9. 9.
    a ship is sailing around on the water
  10. 10.
    a woman with blonde hair and a black shirt is talking
  11. 11.
    a buffalo is attacking a man
  12. 12.
    a band is playing music and people are dancing
  13. 13.
    a child is playing a video game
  14. 14.
    a person is showing how to fold paper
  15. 15.
    a woman is sitting down on a couch in a room
  16. 16.
    a man inside of a car is using his finger to point
  17. 17.
    a man waters his plants
  18. 18.
    the symmetrical cone is japan s most famous symbol
  19. 19.
    an indoor soccer game
  20. 20.
    a japanese monkey bathing in a hot spring with pleasant music
  21. 21.
    some images of motorcycles are being shown on tv
  22. 22.
    someone is serving food in the restaurand
  23. 23.
    this is a competition type show
  24. 24.
    a woman on the news is talking about a story
  25. 25.
    this is a phone review video
  26. 26.
    some fake horses are standing around in a game
  27. 27.
    a person is filming a white car interior seat
  28. 28.
    video of clips from a movie
  29. 29.
    a man with a blue and white shirt is walking around
  30. 30.
    person making something in the kitchen
DEVIL-high
  1. 1.
    A bookshelf collapses loudly, books flying everywhere, creating chaos in the once quiet room.
  2. 2.
    Swift scenes of a sandstorm engulfing a desert oasis, with dunes shifting and palm trees bending in the relentless wind.
  3. 3.
    A chaotic scene of cowboys rounding up cattle during a stampede.
  4. 4.
    Suddenly, a storm hits the city, rain pouring down like a torrent, making rivers on the streets.
  5. 5.
    WWI biplanes in a dogfight with canvas wings ripping, dramatic cloud backdrop, ultra-detailed.
  6. 6.
    In the mountains, a bear erupts from the snow, creating a large cloud of powder.
  7. 7.
    Amidst a thunderstorm, a lightning bolt strikes a bicycle, setting it ablaze with crackling energy and lighting up the dark, rainy street.
  8. 8.
    A single eagle dives extremely fast, snatching a fish from the water.
  9. 9.
    A boat hits a big wave and flips, landing upside down.
  10. 10.
    A car drives through a wall of fire in a daring escape.
  11. 11.
    The cat tore across the living room, jumping over toys and furniture to catch the mouse.
  12. 12.
    A cow jumps over a fence, landing in a pond with a big splash.
  13. 13.
    Two dogs chase each other, suddenly skidding around a sharp corner.
  14. 14.
    A storm sweeps an elephant into a raging river, carrying it away swiftly.
  15. 15.
    Racing the sunset, a giraffe charges across the horizon, shadows stretching long.
  16. 16.
    Against the wind, a lone horse gallops, mane streaming behind.
  17. 17.
    Jumping over a gorge, the motorcycle lands just in time on the other side.
  18. 18.
    A thief sprints away from the scene, with the police in hot pursuit.
  19. 19.
    The ice cracks beneath their feet, making the sheep skid and slide, rushing to solid ground.
  20. 20.
    Lightning strikes as a train blasts its horn, cutting through a stormy night.
  21. 21.
    A truck speeds across the desert, dust clouds swirling behind it.
  22. 22.
    Under a rainbow, a zebra kicks up a spray of water as it crosses a fast-flowing river.
DEVIL-medium
  1. 1.
    London heathrow, united kingdom - 05 12 2019: 4k super-telephoto plane accelerates down hot runway through heat shimmer
  2. 2.
    A cool dj teddy bear with sunglasses on top of turntable with video static
  3. 3.
    Aerial view. cute girl in the coat drive on country road on the bicycle
  4. 4.
    Brown pelican flying flight in fall bay harbor in ecuador
  5. 5.
    Small fishing boat, anchored on a silver ocean, in thailand.
  6. 6.
    a filled yellow school bus with over-sized black wheels drives through a flooded area with red lights on and gets splattered with mud
  7. 7.
    St. petersburg, russia - circa march, 2015: vehicles drive on city ringroad at evening time. st. petersburg ring road is a main route encircling the city
  8. 8.
    cat manages to hang on to dangling object
  9. 9.
    Taking cow milk cheese with fork 4k footage
  10. 10.
    dog passes in and out of view
  11. 11.
    1930s: elephant roars, man shoots at elephant. elephants walk through jungle. man tries to fire gun, throws gun on ground, runs away.
  12. 12.
    the baby giraffe is zoomed in on and then camera shakes
  13. 13.
    Cowboys drive group of horses at farming enterprise.
  14. 14.
    4k couple watching film or tv at home & jumping with shock at the action
  15. 15.
    contestants are reading themselves to start a mini-motorbike race
  16. 16.
    Macao beach with stone mountains aerial view from drone. travel destination. summer vacation. dominican republic
  17. 17.
    Male boxer resting and sweating after boxing training
  18. 18.
    Wild tulips in a meadow on background sky. sunrise. bonfire. a quiet spring morning in the steppe.
  19. 19.
    Sheep eating grass in punata and potosi, bolivia.
  20. 20.
    Bodo arctic town norway - ca july 2018: train station building and rails tilt up
  21. 21.
    a woman is describing different sets of tubes and hoses in the back of a white pick up truck which is parked on the side of a street with cars going by in the background
  22. 22.
    Istanbul, turkey - october 2018: commuters inside istanbul metro wagon travelling towards taksim station
DEVIL-static
  1. 1.
    airplane with red body is shown for first time.
  2. 2.
    a man holds up a stuffed bear.
  3. 3.
    when you can see the first view of the full bike
  4. 4.
    second bird lands on feedersecond bird lands on feeder
  5. 5.
    a red boat is first seen.
  6. 6.
    Tourist bus station 3d realistic footage. public transport front view animation. vehicles on modern urban highway bridge background. passengers transportation parking. city bus stop video
  7. 7.
    black car is under the blue sign.
  8. 8.
    cat looks at the camera
  9. 9.
    dog puts paws together
  10. 10.
    a white horse standing beside red colored wearing girl dress standing with stick bending down knee displaying on screen
  11. 11.
    Blurred conference room with audience - 4k video
  12. 12.
    first time we see orange branch to the right
  13. 13.
    A woman and a man. holding a gift.
  14. 14.
    A tranquil tableau of the old red barn stood weathered and iconic against the backdrop of the countryside
  15. 15.
    black numbers 1758 at bottomof train
  16. 16.
    a large white box truck travels through water is followed by two other trucks and ascends a gray road through mountains
  17. 17.
    view of big city from balconyview of big city from balcony
DEVIL-very-high
  1. 1.
    Classical style of a horse partaking in an ancient chariot race, scenes switching quickly from cheering crowds to close-ups of intense wheel clashes.
  2. 2.
    High-speed shots of a volcanic eruption engulfing a tropical island, with lava fountains spewing molten rock and the environment transforming from idyllic paradise to hellish landscape of ash and fire.
  3. 3.
    A thrilling scene of rural mountain biking extreme sports, starting from the early morning cycling adventure, transitioning to the intense chase through fields and forests, and ending with cheers and celebrations under the sunset.
  4. 4.
    Neon-lit streets pulse with energy as vehicles engage in a high-octane pursuit, transitioning seamlessly from chaos to calculated evasion. Against the backdrop of a setting sun, the chase intensifies, each turn a heartbeat away from capture.
  5. 5.
    A fighter jet dodging rapid anti-air gunfire, quick maneuvers, tracer rounds visible.
  6. 6.
    Bear escaping a collapsing cave, rocks tumbling, dust rising, ((masterpiece)), ((best quality)), 8k, high detailed, ultra-detailed, bear, ((dark rocky textures)), sprinting, (echoing rumble), sudden movement
  7. 7.
    A courier on a bike weaves through traffic at breakneck speed, narrowly avoiding cars and pedestrians in a rush to make deliveries on time.
  8. 8.
    A hummingbird rapidly darting between vibrant flowers in a lush garden, with quick cuts to various close-up shots showcasing its rapid wing movement and agility.
  9. 9.
    Venetian gondola chase scene, narrow canals, historic buildings, urgent escape, ((masterpiece)), ((best quality)), 8k, high detailed, ultra-detailed, gondola, ((twisting canals)), (ancient architecture), (urgent paddling), cinematic chase.
  10. 10.
    Futuristic sports cars racing on a vertical loop track against a sci-fi cityscape, cars defying gravity, ((speed trails)), (dizzying heights), (spectacular crashes), the thrill of cutting-edge technology.
  11. 11.
    Cat rapidly zigzagging across a rooftop, avoiding swooping birds under a stormy sky.
  12. 12.
    A sequence of a cow performing acrobatic stunts over a series of colorful, abstract platforms that morph shapes.
  13. 13.
    The dog bursts through a thicket, darting from a foggy forest to a steep hillside, rocks crumbling under its paws as it charges towards a roaring river below.
  14. 14.
    Thundering across a vast desert plain, the elephants race over dunes and dodge sandstorms, before swiftly traversing through a rocky canyon, bounding over boulders and leaping across narrow ravines.
  15. 15.
    A giraffe navigating a city during a robot uprising, with quick cuts showing chaotic battles, explosions, and futuristic technology in a high-stakes escape scenario.
  16. 16.
    A horse leading a wild stampede across a stormy beach with waves crashing, depicted with swift, sweeping camera moves, cinematic composition.
  17. 17.
    Intense motorcycle escape from a volcanic eruption, with transitions from lava-filled landscapes to ash-clouded skies.
  18. 18.
    A futuristic robot uprising, ((lasers firing)), metallic drones, explosions, debris, ((screaming civilians)), dystopian cityscape.
  19. 19.
    Sheeps engaging in a high-speed pursuit through a cyberpunk city, the scene rapidly transitioning between neon-lit streets, bustling marketplaces, and towering skyscrapers.
  20. 20.
    A pulse-pounding sequence of a train barreling through a treacherous storm, the scene transitioning between lightning-lit skies and torrents of rain to flooded tracks and collapsing bridges.
  21. 21.
    A truck rushing away from a treacherous mountain pass during a blizzard, with sudden avalanches and rockslides adding to the danger.
  22. 22.
    A zebra sprinting across the busy lanes of Times Square in New York City, with scene transitions occurring quickly as it moves from iconic billboards to bustling sidewalks filled with tourists.
MovieGen
  1. 1.
    A green monster made of plants walks through an airport.
  2. 2.
    A marble goes through a glass cup, breaking it into pieces.
  3. 3.
    A droplet of water falling onto a hot surface, instantly evaporating into a wisp of steam that swirls gracefully into the air.
  4. 4.
    An old man wearing a green dress and a sun hat taking a pleasant stroll in Johannesburg South Africa during a beautiful sunset.
  5. 5.
    A person on a hoverboard colliding with a wall, the board stopping abruptly.
  6. 6.
    A toy robot wearing blue jeans and a white t-shirt taking a pleasant stroll in Johannesburg South Africa during a winter storm.
  7. 7.
    In a marathon race, a female athlete gradually sprints ahead of the male athletes.
  8. 8.
    A teenager eating a slice of pizza, cheese stretching as they pull it away.
  9. 9.
    A man in a suit fights monsters.
  10. 10.
    A dog made of ice melts completely in a hot summer day.
  11. 11.
    A truck right alongside a flowing river, capturing the movement of the water and the surrounding forest.
  12. 12.
    A group of skateboarders perform tricks on ramps and rails at a skate park, showcasing their skills.
  13. 13.
    A hot air balloon descending back to the ground.
  14. 14.
    Chef chopping onions in the kitchen for the preparation of the dish.
  15. 15.
    Zoom in shot to the face of a young woman sitting on a bench in the middle of an empty school gym.
  16. 16.
    The couple runs hand in hand to release a sky lantern, then watches it drift upward into the night sky, carried by the wind with the stars shining above.
  17. 17.
    Aerial view shot of a cloaked figure elevating in the sky between skyscrapers.
  18. 18.
    A softball player sliding safely into second base.
  19. 19.
    A giraffe in a lifeguard outfit, sitting atop a high chair and watching over a crowded pool.
  20. 20.
    A speed skater accelerating during a short track race.
Appendix G Detailed Analysis on Reward Function for Perceptual Video Quality
Figure 9 shows that different metrics in reward functions for perceptual video quality often exhibit negative or weak correlations. For example, dynamic degree tends to be negatively correlated with many other metrics, indicating that optimizing exclusively for one metric can either reduce motion dynamics or undermine temporal consistency and aesthetic quality. These findings underscore the need to balance potentially conflicting reward functions rather than prioritizing any single one in isolation, emphasizing the importance of a carefully calibrated approach to evaluating generated videos.
Figure 9: Correlation between reward functions for perceptual video quality.
Appendix H Prompt of AI Feedback
Prompt for AI Feedback from VLMs You are a helpful assistant that evaluates the quality of a generated video from a textual prompt.
Compare the text prompt and generated video and evaluate the quality (visual quality, proper dynamics, etc…) of the video.
First explain the reasoning, then present the final assessment. Start the reasoning with ’Reasoning: ’.
After explaining the reasoning, present the final assessment with ’Assessment: ’.
Your final ’Assessment’ should be a single-number score from 1 to 10, not as a fraction.
When evaluating, consider the following points:
- Visual Quality: Evaluate the clearness, resolution, brightness, aesthetic appeal of the video.
- Dynamics: Evaluate whether the video demonstrates appropriate dynamics, ensuring it avoids excessive movement in situations meant to be static or insufficient movement in situations intended to be dynamic.
- Smoothness, Consistency, and Naturalness: Assess the smoothness, consistency, and naturalness of shape and motion for objects, animals, and humans.
- Contents: Evaluate whether the video content aligns with the given text prompt.
Textual Prompt: {instruction}
Video: {video_file}
Appendix I Further Results for Calibrating Reward to Preference Feedback
I.1 Basic Prompts
Figure 12 and Figure 13 shows the two dimensional histogram and correlation between reward function and AI feedback from Gemini [14] and GPT-4o [13], and Figure 10 represent the coefficient of calibrated reward designed for GPT-4o. The relative weighting of the dynamic degree shifts in accordance with the prompt-specified dynamics, reflecting a pattern similar to that observed in the reward calibration approach with Gemini (DEVIL-high, DEVIL-medium, and DEVIL-static denote prompt sets characterized by high, medium, and low levels of dynamics, respectively). GPT-4o exhibits a stronger inclination toward dynamics than Gemini.
Figure 10: Coefficients of calibrated reward with GPT-4o.
I.2 Challenging Prompts
This section describes the reward calibration procedure and results for two challenging prompt sets, DEVIL-very-high and MovieGen, which were introduced to evaluate our method with larger T2V models. Following the methodology from for reward calibration with Latte (see Section 4), we generated 64 videos per prompt using Wan 2.1-1.3B [20]. Consistent with observations in Appendix I.1, using solely Text-Video Consistency is insufficient to fully capture AI feedback from Gemini (Figure 14). We choose the combination of weights $w_{i}$ to maximize correlation with Gemini’s evaluations. The coefficients of calibrated weights are shown in Figure 11.
Figure 11: Coefficients of calibrated reward with Gemini.
Figure 12: 2D-histogram and correlation between reward function and AI feedback from Gemini.
Figure 13: 2D-histogram and correlation between reward function and AI feedback from GPT-4o.
Figure 14: 2D-histogram and correlation between reward function and AI feedback from Gemini for challenging prompt sets, DEVIL-very-high and MovieGen.
Appendix J Correlation between VLM and Human Evaluation
As mentioned in the prior research [44, 45], VLMs such as Gemini and GPT-4o have been shown to exhibit a high correlation with human evaluations compared to other existing metrics. As an experiment, we measured the correlation between the AI feedback from these VLMs and human labels in the TVGE [72] dataset. As shown in Figure 15, Gemini achieved a correlation of 0.49, and GPT-4o achieved 0.51. Consequently, optimizing for these VLM rewards is a valid way to improve human-perceived quality, rather than merely “gaming” the metrics.
Figure 15: Correlation between VLM outputs and human labels in the TVGE dataset.
Appendix K Qualitative Evaluation of Calibrated Reward
We provide example videos among individual rewards and VLM calibrated rewards in Figure 16. Videos selected solely on a single metric can over-optimize one aspect while neglecting others, whereas those chosen via VLM-calibrated rewards exhibit a more balanced quality. For instance, videos chosen solely based on temporal consistency (i.e., subject consistency and motion smoothness) or frame-by-frame quality (i.e., aesthetic quality, imaging quality) tend to lack dynamic movement, whereas those selected based on dynamic degree often lose temporal consistency. Evaluations relying on a single metric also fail to reflect the given prompt in some cases. Text-video consistency, which often exhibits a high correlation with VLM-based evaluation among individual metrics (Figure 3), is relatively effective in capturing the overall quality of a video. However, it may overlook certain aspects, such as frame-wise artifacts. In contrast, videos selected using VLM-calibrated rewards exhibit a more balanced overall quality.
Figure 16: We select the video with the highest reward out of 64 randomly generated candidates for each prompt, drawn from DEVIL-high, DEVIL-medium, DEVIL-static, and MSRVTT-test (arranged from top to bottom). Videos chosen using VLM-calibrated rewards achieve a more balanced quality compared to those relying on any single metric. For instance, when subject consistency, motion smoothness, or aesthetic quality serves as the sole selection criterion, the resulting videos often lack dynamic movement, whereas prioritizing dynamic degree can compromise temporal consistency. Moreover, single-metric evaluations may occasionally fail to align with the intended prompt.
Appendix L Further Results for Diffusion Latent Beam Search
L.1 Scaling Search Budget with GPT-4o Calibrated Reward
We measure the performance in terms of a combinational reward calibrated to GPT-4o (Figure 17). DLBS improves all the calibrated rewards the best as the search budget $KB$ increases (especially $KB=16,32$), while BoN and GS, in some cases, eventually slow down or saturate the performance. Notably, an LA estimator with a small search budget ($KB=8,T^{\prime}=6$) is comparable to or even outperforms DLBS ($KB=32$).
Figure 17: Inference-time search on reward calibrated to GPT-4o.
L.2 Scaling Search Budget to Larger Regimes
Figure 18 and Figure 19 show the performance of inference-time search on DEVIL-medium and MSRVTT-test that includes the results with $KB=64$. We can observe that the increasing trends still continue.
Figure 18: Inference-time search on reward calibrated to Gemini including $KB=64$.
Figure 19: Inference-time search on reward calibrated to Gemini including $KB=64$.
L.3 Full Results for Scaling Trend of DLBS
Figure 20 demonstrates the scaling trend of DLBS, proportional to the search budget, under various choices of $K$. The results show that there is an optimal balance between the number of latent $K$ and the number of beams $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we have a peak around $K=4,8,16$, which is about 25–50% of the budget. This implies that balancing possession and exploration of diffusion latents in DLBS helps search for the best outputs robustly.
Figure 20: DLBS can improve the performance in any prompts or reward, as we increase the search budget $KB\in\{1,2,4,8,16,32\}$. In addition, we can see an optimal balance between the number of latent $K$ and the number of beam $B$ under the same budget. For instance, as we increase the budget to $KB=16$, we have a peak around $K=4$ or $K=8$, which is about 25–50% of the budget.
L.4 Further Analysis on Lookahead Estimator
Figure 21 (Left) demonstrates that increasing the number of reward estimation steps $T^{\prime}$ in the LA estimator leads to improved reward prediction performance for $\mathbf{z}_{t}$ during the denoising process. This finding suggests that extending the LA steps enables a more effective search based on accurate reward predictions, particularly in the early stages of the denoising. As shown in Figure 21 (Right), enlarging the look-ahead horizon increases the reward gain to $T^{\prime}=6$; beyond this point, e.g., $T^{\prime}=20$ offers no significant benefit while multiplying the calculation. Accordingly, we fix $T^{\prime}=6$ in the main experiments, as it captures nearly all the attainable gains at minimal cost.
Figure 21: (Left) Comparison of the reward estimation error for different LA steps $\mathbf{T^{\prime}}$. We evaluate the reward predicted by the LA estimator, which projects $\mathbf{z}_{t}$ to $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$ in $T^{\prime}$ steps (Algorithm 2) and computes $r^{\prime}(\tilde{\mathbf{z}}_{0|\tilde{t}(0)})$, against the actual reward obtained by projecting $\mathbf{z}_{t}$ to $\mathbf{z}_{0}$ in $t$ steps using a DDIM sampler ($\eta=0.0$) and evaluating $r^{\prime}(\mathbf{z}_{0})$. (Right) Impact of $T^{\prime}$ on search performance. Reward improves rapidly up to $T^{\prime}=6$ but saturates thereafter; using $T^{\prime}=20$ offers no measurable gain. These results show that a modest $T^{\prime}$ is sufficient in practice.
L.5 Ablation Study for Diffusion Steps
Scaling Diffusion Steps  Figure 22 (Left) shows the performance when increasing the number of denoising steps $T$. Since DDIM exhibits fast convergence [42], BoN with larger $T$ does not improve the reward much. DLBS improves the performance when scaling denoising steps to $T=200$ more than BoN, which implies that DLBS benefits from larger computation and the other axes. However, as Figure 6 (Left) indicates, these gains are smaller than those obtained by widening the beam budget $KB$ or leveraging the LA estimator.
Range of Diffusion Steps for Search  We investigate which range of diffusion steps DLBS should be applied to for effective search. Unlike Kim et al. [15], which apply GS only during the initial 5–10 steps, our results in Figure 22 (Right) show that applying DLBS throughout all steps leads to substantially better performance. This suggests that applying DLBS entirely is more effective than focusing on early stage.
Figure 22: (Left) Scaling the denoising steps $T$. (Right) Range of denoising steps $t\in[50,0]$ to apply search methods. While Kim et al. [15] applies GS in the first 5–10 steps, DLBS over the entire diffusion steps yields the largest improvement.
L.6 Ablation Study for $\eta$ in DDIM scheduler
Figure 23 illustrates how varying the value of $\eta$ in DDIM influences search performance. Here, $\eta$ controls the degree of randomness in the DDIM scheduler: $\eta=0.0$ corresponds to the deterministic version of DDIM, while $\eta=1.0$ is equivalent to DDPM. As $\eta$ decreases below $1.0$, performance in terms of the final reward diminishes, presumably because lowering the randomness in the sampling process narrows the scope of exploration.
Figure 23: Comparison among different $\eta$ in DDIM sampler.
L.7 Alignment–Diversity Tradeoff
Alignment for diffusion models can steer desirable outputs, but it is said that the diversity of generated samples or the performance of original models often degrade [10, 37]. While inference-time search does not change or degrade the model itself, we here compare the diversity of samples among BoN, GS, and DLBS. We measure the sample diversity as the mean pairwise distance of ViCLIP [51] embeddings to quantify the diversity in videos, inspired by the approach for evaluating diversity in images [58]. Specifically, given $N$ generated video samples, we first extract ViCLIP embeddings $\mathbf{v}^{\text{video},(i)}$ for each sample $i$. The pairwise diversity score is then computed as the mean pairwise distance:
$D_{\text{video-diversity}}=\frac{1}{N(N-1)}\sum_{i\neq j}\left(1-\langle\mathbf{v}^{\text{video},(i)},\mathbf{v}^{\text{video},(j)}\rangle\right).$ (15)
[…]
Figure 24 reveals that DLBS achieves high performance with higher diversity than BoN or GS. This exhibits a benefit from the wider possession and exploration of diffusion latents in DLBS.
Figure 24: Alignment–diversity tradeoff ($KB=32$), which is measured by the mean pairwise distance of ViCLIP embeddings. DLBS achieves high performance maintaining higher diversity.
L.8 Comparison with Other Inference-Time Search Methods
Concurrently with our work, inference-time search based on sequential Monte Carlo (SMC) has been proposed. We include a comparison with FK Steering [41] in Figure 25, where the resampling mechanism in the SMC-based methods does not occur frequently enough, preventing them from surpassing BoN performance, in our text-to-video experiments.
Figure 25: Comparison with FK steering [41].
L.9 DLBS with Larger Text-to-Video Models
We have tested our method on VideoCrafter2 [3] (1.9B) and CogVideoX-5B, 2B [19] and Wan 2.1-14B, 1.3B [20]. Our experiments confirm that the proposed approach yields significant improvements, indicating its effectiveness can be observed in larger video generation models.
Figure 26: Inference-time search with CogVideoX-5B [19].
Figure 27: Inference-time search with CogVideoX-2B [19].
Figure 28: Inference-time search with Wan 2.1-14B [20].
Figure 29: Inference-time search with Wan 2.1-1.3B [20].
Figure 30: Search with VideoCrafter2 [3].
Figure 31: Search with CogVideoX-5B [19].
L.10 Further Results in AI and Human Evaluation
We show full results of evaluations using VideoScore [54], a metric trained on human judgments that evaluates videos at 8 fps across five dimensions and outputs scores ranging from 1.0 to 4.0 (see Figure 32). Under this metric, videos generated with the proposed search strategy consistently outperformed those generated without search.
Figure 32: Comparison of VideoScore [54]. We compare our approach in the $KB=1$, Best-of-N (BoN), and Diffusion Latent Beam Search (DLBS).
We also show full results of human judgment by three human evaluators (see Figure 33). These experiments confirmed that, regarding human preference (win rate), content generated with our search strategy consistently outperformed content produced without search.
Figure 33: Human evaluation results. We searched videos using the Gemini calibrated reward and asked three human evaluators to compare outputs from GS ($KB=1$) and DLBS-LA. "Win" indicates that the video generated by DLBS-LA was preferred.
L.11 Qualitative Results for DLBS
Qualitative results for DLBS are shown in https://sites.google.com/view/t2v-dlbs.
L.12 DLBS for Image Generation
We adopt PixArt-$\alpha$ [66] as our base text-to-image generation model. For evaluation, we directly reuse the prompt set of 45 common animal categories from prior works [73, 60] to ensure comparability. As a reward model, we employ the LAION aesthetic predictor [49] to assess image quality.
Figure 34: Inference-time search with PixArt-$\alpha$ [66].
[…]
Appendix M Extended Related Works
Aligning Diffusion Models via Finetuning   Alignment by finetuning text-conditioned models has been investigated for image [10] and video [45, 74] generation. Typically, LoRA [75] in a backbone model [76] is finetuned through policy gradient [73, 77, 78], direct preference optimization [79, 44, 80, 38, 81], reward-weighted regression [82], or direct reward gradient [83, 84, 74, 85]. Some train an extra model for better initial noise space [86, 87, 88]. In contrast, we focus on the search over the denoising process at inference time, which does not require any model updates and may not degrade the original performance.
Evaluation of Text-to-Video Generation   While there are several conventional metrics for video generation (or the one repurposed from image generation) such as SSIM [89], IS [90], LPIPS [91], or FVD [92], those are not always suitable to evaluate how the quality of contents in video is, which is much more emphasized in text-to-video generation [72]. It has been a long-standing challenge to comprehensively and semantically evaluate the dynamics of contents or physical commonsense in generated videos [9, 52]. To deal with that, VBench [12] has recently been proposed as a suite of holistic evaluations for text-to-video generation to reflect the perceptual aspect of the quality, such as consistency, smoothness, aesthetics of contents, or text–video alignment. Moreover, inspired by the success in LLMs [93, 94, 95], we could leverage VLMs, which become more capable these days, as a proxy of human evaluation of the contents [74]; by finetuning CLIP-based models [54, 68, 96], or prompting GPT-4o [13] or Gemini [14]. Our paper adopts AI feedback from VLMs as an alternative to human rater, and proposes a recipe to calibrate a reward to other sources of feedback (such as AI or human feedback), by considering a linear combination of fine-grained metrics.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our abstract and introduction appropriately include the claims made in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See section 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe the details of experiments in Section 5 and other necessary information in Appendix B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our experiments are based on the open-source dataset [52, 53, 55]. Our experimental code is shown in https://anonymous.4open.science/r/T2V-Diffusion-Search-537B.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe the details of experiments in Section 5 and other necessary information in Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report the statistical significance of the Pearson correlation coefficient in Figure 3 and Appendix I.1. We also report the reward values averaged over multiple prompts to reduce the variance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Appendix B.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We believe our research conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not include new datasets or pre-trained models that pose a risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have appropriately cited the papers of existing assets we used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: he paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [No]
  69. Justification: We only recruited participants for user experiments to validate the effectiveness of our model, where they were asked to choose from generated videos. No human participants were involved in the dataset construction or model training process.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our experiment solely involves measurement and does not entail behavioral manipulation; therefore, we did not apply for IRB approval.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We use LLMs only for writing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.